Can you please help me solve the following equation?
-2(3b+1)+(7b-4)=0
Thank you so much and God bless you!

Answer by RAY100(1637)

-2(3b+1) + (7b-4) =0
-6b-2+7b-4=0
b-6=0
add 6 to both sides of eqn
b=6
check
-2(3(6)+1) +(7(6)-4) =0
-2*19+38 =0 ok